How many grams of ammonia can be made by reacting 100 g of N2 with 20 g of ?

What is the limiting reactant?

How much of the excess reactant will be left over? N2(g) + 3H2(g) → 2NH3(g)


1
Expert's answer
2021-09-21T01:55:20-0400

Molar mass of N2 = 28.0134

100/28.0134 = 3.57moles

Molar ratio of N2:NH3 = 1:2

3.57 × 2 = 7.14moles of NH3

Molar mass of H2 = 2.016

20/2.016 = 9.92moles

Mole ratio of H2:NH3 = 3:2

9.92×2/3 = 6.613 moles

= 6.613+7.14 = 13.753 moles

Molar mass of NH3 = 17.031

17.031×13.753 = 234.233g

N2 is the limiting reactant